OpenStudy (anonymous):

write the equation in the slope-intercept form 11x + 9y = 5

OpenStudy (anonymous):

Do you know what the slope-intercept equation looks like?

OpenStudy (anonymous):

Solve the equation for y...the slope intercept form of an equation is in the form y=mx+b

OpenStudy (anonymous):

11x + 9y = 5 11x - 11x + 9y = 5 -11x 9y = 5 -11x 9y/9 = ( 5 -11x)/9 y=5/9-11x/9

OpenStudy (anonymous):

Why just give her the answer? How have you helped her?

OpenStudy (anonymous):

well slope intercept form is y=mx+b another method to so this is to find the slope from the equation using -a/b standard form is ax+by=c in this case the slope would be -11/9 or m=-11/9 the 2nd part of the equation is the y-intercept which we can find if we substitute x=0 11x+9y= 5 11(0)+9y=5 9y=5 y=5/9 so the y intercept is 5/9 or b=5/9 now that we have both portions needed we input them into y=mx+b y=-11/9x+5/9